Joe is going to put wallpaper on all walls of the bedroom. here is the formula

R=2L+2w
________
2.5
R = wallpapers needed L=Length of room(m)W=width of room(m)
The bedroom has a length of 5.6m, A width of 4.8
How many rolls of wallpaper does joe need to buy, explain why?

Answer by fractalier(6550)

I would like to help you on this one (you posted it twice) but since wallpaper goes on walls that have heights, and you did not supply the heights of the walls, I cannot help you.
If all you want is
R = 2L + 2W = 2(5.6) + 2(4.8) = 11.2 + 9.6 = 20.8 m
Your use of 2.5 isn't clear...did you want to divide the total above by 2.5?
If so, 20.8/2.5 = 8.32 rolls...he would need to buy 9 rolls...